OpenStudy (anonymous):

y squared = - 7 solve for Y

OpenStudy (anonymous):

If this is in the real number system then there is no solution since \[y^2= positive\] for real numbers For Complex numbers it would be \[y=\pm \sqrt{(-1)(7)}=\pm i\sqrt{7} \]

OpenStudy (anonymous):

Okay, sorry, my daughter didn't give me the full equation. It is -(ysquared-2) = ysquared -5-2ysquared

OpenStudy (anonymous):

\[-(y^2-2)=y^2-5-2y^2\] ?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

The answer should then be no solution since \[-y^2 +2 = y^2-2y^2 -5 \to -y^2=-y^2-7\to y^2 = y^2+7\] There is no value of y that when you square y^2 will equal itself squared plus a number y^2 +7.